# Language Models, Graph Searching, and Supervision Adulteration: When More Supervision is Less and How to Make More More

### v1

@@ 1 Introduction
The path-star task is a seemingly simple minimal graph search task intended to exhibit a flaw in the standard next-token prediction paradigm used to train decoder-only autoregressive LMs via teacher-forcing (TF) (Bachmann and Nagarajan, 2024).
Each graph is star-shaped with $D$ arms rooted at a single start node, $s$. The LM is given the complete graph (as a shuffled edge list) and a query, $(s,\,t)$, where $t$ is a target node that ends an arm. The task is to generate the arm with $t$ from $s$ to $t$ (Fig 1). This requires the LM to choose an arm by initially generating one of the $D$ leading nodes adjacent to $s$ with the rest of the arm being dictated by following edges. Thus the difficulty lies in choosing the correct leading node, $l_{t}$, which necessitates planing and reconstruction of the correct arm from $t$ to $l_{t}$.
Training via TF conditions the LM on prior ground-truth tokens. This induces learning an undesired shortcut, the Clever Hans Cheat (CHC), which allows for trivial prediction of all non-leading nodes via a single edge look-up given the preceding node (given via TF). Thus all the sequential supervision is absorbed into learning the CHC except a single target token, $l_{t}$, which becomes the sole support for learning the required arm reconstruction subtask. [^1] Why this itself is hard is an open question, see Sec. 2.2. As a result, LMs fail to generate the correct arm above the random baseline of $1/D$ chance (Bachmann and Nagarajan, 2024). While it has been shown that decoder-only LMs can express the task (Frydenlund, 2024), it remains an open question if decoder-only language models trained via teacher-forcing can learn the task.
@@ 1.1 Significance of the Failure
LMs are the ubiquitous model for NLP tasks (Brown et al., 2020), as well as for reasoning tasks (Bubeck et al., 2023). These tasks often require planning, which LMs struggle with (Valmeekam et al., 2023b; Kambhampati et al., 2024, Ap. B.1.1).
@@
This motivated using alternative models. Bachmann and Nagarajan (2024) used a ‘teacher-less’ model which foregoes TF by conditioning on fully masked input (Monea et al., 2023). Frydenlund (2024) generalized this to non- and iterative-autoregressive models and demonstrated learnability differences between models, where an encoder-only LM could solve the task (on small graphs).
Saparov et al. (2025) showed positive results on path-star graphs with encoder-only models and on more general graph topologies with both encoder- and decoder-only models. They did not try path-star graphs with decoder-only LMs. They found that the topology is critical to generalization but that learning does not scale with graph size (and using scratchpads which perform a depth-first-search did not resolve this issue), leading to the claim that ‘transformers struggle to learn to search’. They also found that each node learns to store all other reachable nodes, suggesting the models learn an unscalable representation to solve search tasks.
Yin et al. (2024) and Hu et al. (2025a) both proposed novel model architectures on the perceived deficiency of decoder-only models in solving the path-star task. Yin et al. (2024) trained an auxiliary autoencoder to form planning latent-states that encode future tokens. Then they trained an LM with special planning tokens regressed against the autoencoder’s planning latent-states. These planning tokens are provided as input during inference, thus providing information about future tokens. Hu et al. (2025a) introduced a model trained on both forward and backward contexts using two separate forward and backward encoders. These encoders then condition two separate forward and backward distributions. During inference, the backward encoder and decoder are unused, and the forward decoder conditions on a set of blank states instead of those produced by the backward encoder.
Wu et al. (2024c) put forth a related argument that next-token prediction is potentially problematic for planning tasks due to the cross-entropy loss leading to spurious correlations (see Appx. B.2).
@@ 2 Task, Data, and Tokenization
Each graph, $G$, has $D$ arms of the same length $M$ (inclusive of $s$) and is constructed by sampling nodes from a set of possible nodes, $V$, without replacement, making the graph size $|G|=D(M-1)+1$. The edges are determined by this sample order. Thus all nodes are unique and semanticless as they only relate via randomly sampled edges.
The task is tokenized as a sequence consisting of the query, $Q=(s,\,t)$, with start- and end-of-query markers (‘/ s t ?’). Each edge, $(u,\,v)$, is followed by the end-of-edge marker (‘u v |’). See Fig. 2. The entire graph is provided to the language model as a series of edges which are randomly shuffled. This destroys any higher-order structural information about $G$, meaning that the task must be solved via planning and edge-following. The source-side input into the model is $Q$ followed by $G$ and the end-of-graph marker (‘=’). We place $Q$ before $G$ as it is better for decoder-only models (Frydenlund, 2024). Let $R_{t}=x_{1},\,\dots,\,x_{M}$ be the series of nodes from $s$ to $t$ forming the target arm and the sequential target-side supervision.
Each experiment uses a model trained from scratch on graphs with static $D$ and $M$, i.e. different sized graphs are not mixed during training (except in Sec. 3.6). We avoid uncontrollable biases from natural data by not using pretrained models.
@@ 2.1 Supervision Adulteration
Consider the various ways the task is supervised. In a supervised learning framework, we generally regard the target labels as ‘the supervision’ as they can be human-annotated. With the next-token prediction paradigm, we forego human-annotation by using a self-supervised rule for generating the targets. In both cases, the targets are a function of the input and thus the choice of input is just as much a form of supervision as the targets themselves. Under this view, the model is provided with three types of supervision during training: the target-side labels, target-side inputs, and source-side inputs.
The path-star task (PST) is designed to induce a bad interaction between these three types of supervision under standard training. For a given step $i$, the LM is trained to predict the target-side label $x_{i}\in R_{t}$. However, it will be given $x_{<i}$, including $x_{i-1}$ as target-side input due to TF. This induces learning the trivial single-edge lookup as the edge $(x_{i-1},\,x_{i})$ is provided in the source-side input. [^2] To foreshadow Sec. 3.5, this would not be possible if that edge was not provided in the source-side inputs. Fig. 3(a) illustrates the CHC as a single-edge lookup.
Thus excessive supervision partitions the sequential target-label supervision into supporting two tasks: the desired PST, supported by a single target label, and the undesired single-edge lookup task, supported by the remaining labels. This indicates that the task is not constructed properly to induce learning the PST. We will demonstrate that 1) this bad interaction, and thus the CHC, can be avoided in various ways and 2) avoiding the CHC is not actually critical for learning the task if we consider other ways the task is supervised.
In general, task construction is a form of supervision encompassing multiple design decisions. We considered the target-side above, however, the source-side representation is also supervised. For example, how the $G$ is shuffled matters (edge-wide vs. arm-wise (Frydenlund, 2024) or casual-wise as in Tbl. 1), the decision to place $Q$ after or before $G$ or which tokens to include in the query (Sec. 3.4).
@@ 3 Methods and Experiments
We use decoder-only models with 2 heads, 64 dim. embeddings, 256 dim. feed-forward layers, and learned positional embeddings. We use $L=8$ layers for all experiments. Having $M<L$ allows for the linear graph reconstruction alg. to be learnt. Frydenlund (2024) proved $\mathcal{O}(\log(M))$ layers are sufficient for this in theory. This was empirically demonstrated by Yin et al. (2024); Saparov et al. (2025) who use $L<M$. Sanford et al. (2024b) demonstrated that this $\mathcal{O}(\log(M))$ alg. can be learnt for a related task. We use a learning-rate of $5*10^{-4}$, a batch-size of 1024, and do not use dropout or a scheduler. We train for 100M samples.
@@
We modify the task setting from Bachmann and Nagarajan (2024) by a) placing $Q$ before $G$, b) using an online dataset to avoid overfitting, and c) setting $|V|=|G|$ instead of 100. These changes are immaterial to any conjecture regarding an inability to plan and do not prevent learning the CHC or its apparent effect on the task. We confirm the PST is still unlearnable (even on minimal graphs with only $|G|=9$ nodes) in this setting in Tbl. 1 (full results are in Tbl. 6 in Appx. A.1).
Exp. Desc. $D$ $M$ SR ABB
Edge-Wise 2 5 0% 0%
5 5 0% 0%
2 7 0% 0%
5 7 0% 0%
Causal-Wise 5 5 60% 100%
2 7 100% 100%
3 7 60% 80%
2 9 40% 100%
3 9 0% 40%
2 12 0% 80%
Table 1: Baseline results. We report the Success Rate (SR) where the model predicts $>95$% sequential accuracy over $n=5$ seeded trials and Above-Baseline (ABB) where the model predicts $>(100/D+10)$sequential accuracy. This happens when the model can predict $l_{t}$ above $1/D$ chance. As such, when ABB $>$ SR, it implies that the model has overcome the main challenge of the PST and would have learnt the task had it been provided with more training time in these cases. We report using standard autoregressive generation.
Tbl. 1 shows the PST is learnable with causal-wise shuffling, indicating that the causal constraint accounts for some of the task’s difficulty.
@@ 3.1 Token Masking
A main innovation from these models is that we do not need to employ full masking unlike the ‘teacher-less’ and non-autoregressive models and we can keep the causal parameterization of the model, unlike the iterative- and non-autoregressive models. Importantly, this can be achieved via ubiquitous data-noising methods used with standard TFed training. In particular, we can employ either token dropout/masking (Gal and Ghahramani, 2016; Bowman et al., 2016) or token replacements via scheduled sampling [^4] As we know how the model generates, we skip implementing scheduled sampling and just sample from $V$ instead. (Bengio et al., 2015) (or a mix). Replacement has the benefit of providing an anti-CHC learning signal as the model can not trust edge look-ups, but, it also introduces more complex noise. We use contiguous span sampling (modified to shun consecutive ground-truths) instead of uniform sampling (Joshi et al., 2020).
@@ 3.1.1 Results and Discussion
Exp. Desc. $D$ $M$ SR ABB
Span Token Dropout 2 5 100% 100%
2 7 80% 80%
5 7 40% 40%
3 9 0% 0%
2 12 0% 0%
Span Mixed Token Dropout and Replacement 2 5 80% 80%
2 7 100% 100%
5 7 0% 20%
3 9 80% 80%
2 12 20% 20%
Table 2: Masking results (full Tbl. 8 in Appx. A.2).
Tbl. 2 shows that masking makes the task learnable but struggles as $D$ and $M$ scale. We find minor differences in the two masking types and try mixing them, as they may provide different benefits (token replacement tells the model not to trust single-edge lookups while a masked token prevents these).
Finding that a given method makes the PST learnable but does not scale will be a consistent pattern across methods. Our focus for these methods is, a), showing that the task becomes learnable and, b), explaining why. We conjecture about limitations in scalability in Sec. 5 which were also observed by Saparov et al. (2025, see Appx. B.2).
@@ 3.1.2 Unadulterated Task Decomposition
We show how masking prevents the CHC in Fig. 3. First, consider the CHC in Fig. 3(a) and the needed algorithm for predicting $l_{t}$ in Fig. 3(b). The CHC learns a forward alg. from the prior token, while the needed alg. must work backward from the target query. Figs. 3(c) and 3(d) show how masking induces multi-edge lookups. In Fig. 3(c), when all prior tokens are masked, it induces learning a subset of steps for the needed alg. This provides a deeper explanation for why masking works beyond preventing the CHC; it induces task decomposition. This also explains why unadulterated sequential supervision is important. Decomposition occurs because reconstruction is inherently recursive.
Fig. 3(d) shows having unmasked prior tokens may lead to learning the forward alg. While these may seem similar to a human, they are different algs. and it’s unclear if they mutually support each other in terms of learning to predict $l_{t}$. We tried always masking $l_{t}$ as input, however, this caused worse predictive performance.
@@
(b) Arm reconstruction needed for predicting $l_{t}=$‘12’. Note how the algorithm must work backwards from $t=$‘2’. Steps 4 and 5 are verification steps to match $l_{t}$ being adjacent to $s$. These are not actually learnt due to positional shortcuts.
(c) Predicting ‘6’ with $l_{t}$ masked. This avoids the CHC and induces learning a decomposition or subset of steps needed to predict $l_{t}$. This exact decomposition requires a front-spanning mask which disallows conditioning on any prior ground-truths and explains the (limited) success of the ‘teacher-less’ and non-autoregressive models.
@@ 3.2 Alternative Sequential Distributions
We consider changing the learnt distribution from one over the next token to one over the next tokens. This is done via learning a belief-state, $B$, which is a hidden-state that supports making future predictions via some linear function of $B$ i.e., $P_{B}(x_{i:M}\,|\,f(B=x_{<i}))$ (Hu et al., 2025a).
We present three simple methods for learning this future distribution, $P_{B}$: bag-of-words (BoW), label-smoothing (LS), and ranking. Yin et al. (2024) used a BoW baseline with $R_{t}$ as the bag. This performed nearly as well as their proposed model and solved the task in the majority of cases. [^5] They used pretrained GPT2 models in their experiments which we expect will increase performance and scalability. We exclude prior tokens ($<i$) from the bag so as to only contain future tokens at each step. Note BoW is equivalent to LS with uniform smoothing.
BoW is based on the inductive bias that nodes in $R_{t}$ are more important than nodes in the other arms. We can extend this with another inductive bias which assumes that near-present tokens are more important than far-future tokens i.e. that the order matters. This can be achieved using monotonically decreasing label weights. Thus LS requires hand-crafted weights to form a hand-crafted distribution that the model tries to match. We can avoid this ad-hoc approach via an equivalence between LS and ranking (Frydenlund et al., 2022).

### v3

@@ 1 Introduction
The path-star task is a seemingly simple, minimal graph search task intended to exhibit a flaw in the standard next-token prediction paradigm used to train decoder-only autoregressive LMs via teacher-forcing (TF) (Bachmann and Nagarajan, 2024).
Each graph is star-shaped with $D$ arms rooted at a single start node, $s$. The LM is given the complete graph (as a shuffled edge list) and a query, $(s,\,t)$, where $t$ is a target node that ends an arm. The task is to generate the arm with $t$ from $s$ to $t$ (Fig 1). This requires the LM to choose an arm by initially generating one of the $D$ leading nodes adjacent to $s$, with the rest of the arm being dictated by following edges. Thus, the task’s difficulty lies in choosing the correct leading node, $l_{t}$, necessitating planning and reconstruction of the correct arm from $t$ to $l_{t}$.
Training via TF conditions the LM on prior ground-truth tokens. This induces learning an undesired shortcut, the Clever Hans Cheat (CHC), which allows for trivial prediction of all non-leading nodes via a single edge look-up given the preceding node (given via TF). Thus, all the sequential supervision is absorbed into learning the CHC except for a single target token, $l_{t}$, which becomes the sole support for learning the required arm reconstruction subtask. [^1] Why this itself is hard is an open question, see Sec. 2.2. As a result, LMs fail to generate the correct arm above the random baseline of $1/D$ chance (Bachmann and Nagarajan, 2024). While it has been shown that decoder-only LMs can express the task (Frydenlund, 2024), it remains an open question if decoder-only language models trained via teacher-forcing can learn the task.
@@ 1.1 Significance of the Failure
LMs are the ubiquitous model for NLP tasks (Brown et al., 2020), as well as for reasoning tasks (Bubeck et al., 2023). These tasks often require planning, which LMs struggle with (Valmeekam et al., 2023b; Kambhampati et al., 2024, Ap. C.1.1).
@@
This motivated the use of alternative models. Bachmann and Nagarajan (2024) used a ‘teacher-less’ model which foregoes TF by conditioning on fully masked input (Monea et al., 2023). Frydenlund (2024) generalized this to non- and iterative-autoregressive models and demonstrated learnability differences between models, where an encoder-only LM could solve the task (on small graphs).
Saparov et al. (2025) showed positive results on path-star graphs with encoder-only models and on more general graph topologies with both encoder- and decoder-only models. They did not try path-star graphs with decoder-only LMs. They found that the topology is critical to generalization, but that learning does not scale with graph size (and using scratchpad that performs a depth-first search did not resolve this issue), leading to the claim that ‘transformers struggle to learn to search’.
Yin et al. (2024) and Hu et al. (2025a) both proposed novel model architectures on the perceived deficiency of decoder-only models in solving the path-star task. Yin et al. (2024) trained an auxiliary autoencoder to form planning latent-states that encode future tokens and then trained an LM which regressed against these to learn special planning tokens. Hu et al. (2025a) introduced a model trained on both forward and backward contexts using two separate forward and backward encoders. Ahn et al. (2025) continued with this motivation but tried to minimize the required architecture changes.
Wu et al. (2024c) put forth a related argument that next-token prediction is potentially problematic for planning tasks due to the cross-entropy loss leading to spurious correlations (see Appx. C.2).
1.2 Disproving Prior Claim and Conjecture
The claim that failure to learn the task demonstrates an insufficiency of the standard paradigm is empirically supported. Hence, showing the task to be learnable with standard methods would refute this claim. Bachmann and Nagarajan (2024) also conjectured that this failure will apply to more complex planning tasks. They gave story generation as an example, but without any empirical testing.
This work will show that this claim and conjecture are unfounded, or at least extremely limited, by demonstrating that the task is fragile to minor modifications which make the task learnable via standard methods. We show that the CHC is not the sole cause of the task’s difficulty, and that preventing it is unnecessary for learning the task. We instead show that subtask decomposition is necessary for learning the task, and explain that the failure is caused by excessive or adulterated supervision that prevents this decomposition.
Saparov et al. (2025) showed strong success on the task with encoder-only models and on a similar-looking task with decoder-only models. This success came without discussion with Bachmann and Nagarajan (2024). We will explain this gap as also being attributable to subtle changes in the task that induce subtask decomposition. Like Saparov et al. (2025), we find the task becomes harder to learn with scale. However, it is unnecessary to show scalability to larger graphs to refute prior claims.
@@ 2 Task, Data, and Tokenization
Each graph, $G$, has $D$ arms of the same length $M$ (inclusive of $s$) and is constructed by sampling nodes from a set of possible nodes, $V$, without replacement, making the graph size $|G|=D(M-1)+1$. The edges are determined by this sample order. Thus, all nodes are unique and semanticless as they only relate via randomly sampled edges.
The task is tokenized as a sequence consisting of the query, $Q=(s,\,t)$, with start- and end-of-query markers (‘/ s t ?’). Each edge, $(u,\,v)$, is followed by the end-of-edge marker (‘u v |’). See Fig. 2. The entire graph is provided to the language model as a series of edges, which are randomly shuffled. This destroys any higher-order structural information about $G$, meaning that the task must be solved via planning and edge-following. The source-side input into the model is $Q$ followed by $G$ and the end-of-graph marker (‘=’). We place $Q$ before $G$ as it is better for decoder-only models (Frydenlund, 2024). Let $R_{t}=x_{1},\,\dots,\,x_{M}$ be the series of nodes from $s$ to $t$ forming the target arm and the sequential target-side supervision.
Each experiment uses a model trained from scratch on graphs with static $D$ and $M$, so different-sized graphs are not mixed during training (except in Sec. 3.6). We avoid uncontrollable biases from natural data by not using pretrained models.
@@ 2.1 Supervision Adulteration
Consider the various ways the task is supervised. In a supervised learning framework, we generally regard the target labels as ‘the supervision’ as they can be human-annotated. With the next-token prediction paradigm, we forego human annotation by using a self-supervised rule for generating the targets. In both cases, the targets are a function of the input and thus the choice of input is just as much a form of supervision as the targets themselves. Under this view, the model is provided with three types of supervision during training: the target-side labels, target-side inputs, and source-side inputs.
The path-star task (PST) is designed to induce a bad interaction between these three types of supervision under standard training. For a given step $i$, the LM is trained to predict the target-side label $x_{i}\in R_{t}$. However, it will be given $x_{<i}$, including $x_{i-1}$ as target-side input due to TF. This induces learning a trivial single-edge lookup as the edge $(x_{i-1},\,x_{i})$ is provided in the source-side input. [^2] To foreshadow Sec. 3.4, this would not be possible if that edge was not provided in the source-side inputs. Fig. 5(a) illustrates the CHC as a single-edge lookup.
Thus excessive supervision partitions the sequential target-label supervision into supporting two tasks: the desired PST, supported by a single target label, and the undesired single-edge lookup task, supported by the remaining labels. This indicates that the task is not constructed properly to induce learning the PST. We will demonstrate that 1) this bad interaction, and thus the CHC, can be avoided in various ways and, 2) avoiding the CHC is not actually critical for learning the task if we consider other ways the task is supervised.
Task construction is a form of supervision encompassing multiple design decisions. We considered the target-side above, however, the source-side representation is also supervised. For example, how the $G$ is shuffled matters (edge-wise vs. arm-wise (Frydenlund, 2024) or causal-wise as in Tbl. 3), the decision to place $Q$ after or before $G$ or which tokens to include in the query (Sec. 3.5).
@@ 3 Methods and Experiments
We use decoder-only models with 2 heads, 64 dim. embeddings, 256 dim. feed-forward layers, and learned positional embeddings. We use $L=8$ layers for most experiments. Having $M<L$ allows for the linear graph reconstruction alg. to be learnt. Frydenlund (2024) proved $\mathcal{O}(\log(M))$ layers are sufficient for this in theory. This was empirically demonstrated by Yin et al. (2024); Saparov et al. (2025) who use $L<M$. Sanford et al. (2024b) demonstrated that this $\mathcal{O}(\log(M))$ alg. can be learnt for a related task. We use a learning-rate of $5\times 10^{-4}$, a batch-size of 1024, and do not use dropout or a scheduler. We train for 100M samples.
@@
We modify the task setting from Bachmann and Nagarajan (2024) by a) placing $Q$ before $G$, b) using an online dataset to avoid overfitting, and c) setting $|V|=|G|$ instead of 100. These changes are immaterial to any conjecture regarding an inability to plan and do not prevent learning the CHC or its apparent effect on the task. We confirm the PST is still unlearnable (even on minimal graphs with only $|G|=9$ nodes) in this setting in Fig. 3 (full results are in Tbl. 1 in Appx. B.1).
Figure 3: Baseline results. We report the Success Rate (SR) where the model predicts $>95$% sequential accuracy over $n=5$ seeded trials and Above-Baseline (ABB) where the model predicts $>(100/D+10)$sequential accuracy. This happens when the model can predict $l_{t}$ above $1/D$ chance. As such, when ABB $>$ SR ($\uparrow$), it implies that the model has overcome the main challenge of the PST and would have learnt the task had it been provided with more training time in these cases. An ‘x’ further indicates no trials learnt the task.
Fig. 3 shows the PST is learnable with causal-wise shuffling, indicating that the causal constraint accounts for some of the task’s difficulty. All further experiments use ‘edge-wise’ shuffling.
In the following sections, we will introduce a method or a slight modification to the task which allows the task to be learnt. In general, these will use standard teacher-forcing and next-token prediction. They will also be orthogonal to each other. Importantly, the success of each method can be explained as avoiding supervision adulteration and so inducing subtask decomposition (one example of this is illustrated in Fig. 5).
@@ 3.1 Token Masking
A main innovation from these models is that we do not need to employ full masking, unlike the ‘teacher-less’ and non-autoregressive models and we can keep the causal parameterization of the model, unlike the iterative- and non-autoregressive models. Importantly, this can be achieved via ubiquitous data-noising methods used with standard TFed training. In particular, we can employ either token dropout/masking (Gal and Ghahramani, 2016; Bowman et al., 2016) or token replacements via scheduled sampling [^4] As we know how the model generates, we skip implementing scheduled sampling and just sample from $V$ instead. (Bengio et al., 2015) (or a mix). Replacement has the benefit of providing an anti-CHC learning signal as the model can not trust edge look-ups, but it introduces more complex noise. We try both uniform sampling of the sequence length and contiguous span sampling to shun consecutive ground-truths (Joshi et al., 2020).
@@ 3.1.1 Results and Discussion
Figure 4: Masking results (full Tbl. 3 in Appx. B.2).
Fig. 4 shows that masking makes the task learnable but struggles as $D$ and $M$ scale. We find minor differences in the two masking types and try mixing them, as they may provide different benefits (token replacement tells the model not to trust single-edge lookups while a masked token prevents these).
Finding that a given method makes the PST learnable but does not scale will be a consistent pattern across methods. Our focus for these methods is, a), showing that the task becomes learnable and, b), explaining why. We conjecture about limitations in scalability in Sec. 5 which were also observed by Saparov et al. (2025, see Appx. C.2).
@@ 3.1.2 Unadulterated Task Decomposition
We show how masking prevents the CHC in Fig. 5. First, consider the CHC in Fig. 5(a) and the needed algorithm for predicting $l_{t}$ in Fig. 5(b). The CHC learns a forward alg. from the prior token, while the required alg. must work backward from the target query. Figs. 5(c) and 5(d) show how masking induces multi-edge lookups. In Fig. 5(c), when all prior tokens are masked, it induces learning a subset of steps for the required alg. This provides a deeper explanation for why masking works beyond preventing the CHC; it induces task decomposition. This also explains why unadulterated sequential supervision is important. Decomposition can occur because arm reconstruction is inherently recursive.
Fig. 5(d) shows that having unmasked prior tokens may lead to learning a forward alg. While these seem similar to a human, they are different algs. and it’s unclear if they mutually support each other in terms of learning to predict $l_{t}$, i.e., does a subtask need to exactly mirror the core task for subtask decomposition to help learning the core task.
@@
(b) Arm reconstruction needed for predicting $l_{t}=$‘12’. Note how the algorithm must work backward from $t=$‘2’. Steps 4 and 5 are verification steps to match $l_{t}$ being adjacent to $s$. These are not actually learnt due to positional shortcuts.
(c) Predicting ‘6’ with $l_{t}$ masked. This avoids the CHC and induces learning a decomposed subtask which mirrors the steps needed to predict $l_{t}$ in the core task while also simplifying the task since it is only a subset of the steps. This exact decomposition requires a front-spanning mask that disallows conditioning on any prior ground-truths and explains the (limited) success of the ‘teacher-less’ model.
@@ 3.2 Alternative Sequential Distributions
We consider learning a distribution over the next tokens instead of the next token, i.e., multi-token prediction. This is done via learning a belief-state, $B$, which is a hidden-state that supports making future predictions via some linear function of $B$ i.e., $P_{B}(x_{i:M}\,|\,f(B=x_{<i}))$ (Hu et al., 2025a).
We present three simple methods for learning this future distribution, $P_{B}$: bag-of-words (BoW), label-smoothing (LS), and ranking. Yin et al. (2024) used a BoW baseline with $R_{t}$ as the bag. This performed nearly as well as their proposed model and solved the task in the majority of cases. [^5] They used pretrained GPT2 models in their experiments. We exclude prior tokens ($<i$) from the bag so as to only contain future tokens at each step. Note BoW is equivalent to LS with uniform smoothing.
BoW is based on the inductive bias that nodes in $R_{t}$ are more important than nodes in the other arms. We can extend this with another inductive bias which assumes that near-present tokens are more important than far-future tokens, i.e., that the order matters. This can be achieved using monotonically decreasing label weights. Thus, LS requires hand-crafted weights to form a hand-crafted distribution that the model tries to match. We can avoid this ad-hoc approach via an equivalence between LS and ranking (Frydenlund et al., 2022).